40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 11 00:07:54 2015, rana, Update, LSC, Lock Loss plot 1107673198.png1107673198.png
    Reply  Wed Feb 11 02:42:05 2015, Jenne, Update, LSC, PRC error signal RF spectra REFL33_25.pdfREFL33_15.pdf
       Reply  Wed Feb 11 03:41:12 2015, Koji, Update, LSC, PRC error signal RF spectra 
          Reply  Wed Feb 11 17:31:11 2015, ericq, Update, LSC, RFPD spectra REFL.pngAS.pngPOP.png2015-02-PDSpectra.zip
             Reply  Wed Feb 11 18:07:42 2015, ericq, Update, LSC, RFPD spectra 
                Reply  Thu Feb 12 01:00:18 2015, rana, Update, LSC, RFPD spectra 
Message ID: 11003     Entry time: Wed Feb 11 17:31:11 2015     In reply to: 11000     Reply to this: 11004
Author: ericq 
Type: Update 
Category: LSC 
Subject: RFPD spectra 

For future reference, I've taken spectra of our various RFPDs while the PRMI was sideband locked on REFL33, using a 20dB RF coupler at the RF input of the demodulator boards. The 20dB coupling loss has been added back in on the plots. Data files are attached in a zip.

Exceptions: 

  • The REFL165 trace was taken at the input of the amplifier that immediately preceeds the demod board. 
  • The 'POPBB' trace was taken with the coupler at the input of the bias tee, that leads to an amplifier, then splitter, then the 110 and 22 demod boards. 

I also completely removed the cabling for REFLDC -> CM board, since it doesn't look like we plan on using it anytime in the immediate future. 

Attachment 1: REFL.png  81 kB  | Hide | Hide all
REFL.png
Attachment 2: AS.png  59 kB  | Hide | Hide all
AS.png
Attachment 3: POP.png  99 kB  | Hide | Hide all
POP.png
Attachment 4: 2015-02-PDSpectra.zip  283 kB
ELOG V3.1.3-